# Observation of entanglement in a cold atom analog of cosmological preheating

### v1

@@ Abstract
FIG. 3: Plot of the mean detected population (a) and the cross-correlator (b) as a function of the hold time $\Delta t$. Data is shown for two different modulation depths (orange circles 18%, green squares 25%). The inset shows the normalized variance. It is seen that at long times one cannot infer the existence of any entanglement. Error bars denote one standard deviation uncertainty and are computed using a bootstrap analysis.
Time evolution – We investigate the time evolution by fixing the modulation amplitude and varying the hold time. The results are shown in Fig. 3 for amplitudes of 18% (orange circles) and 25% (green squares). The mean quasiparticle population in the excited modes, shown in Fig. 3(a), exhibits exponential growth, consistent with expectations from parametric resonance [54] and our two-mode squeezed thermal state model  [^4] The growth will be discussed in more detail in a future publication., see Eq. (5). The corresponding evolution of the normalized two-body correlator $g^{(2)}_{\pm}$ is shown in Fig. 3(b) and the normalized variance between the two modes is shown in the inset of the figure. As the hold time $\Delta t$ increases, $g^{(2)}_{\pm}$ gradually approaches the value 2. This behavior is easily understood: although the anomalous correlation $|\langle\hat{a}_{+}\hat{a}_{-}\rangle|$ increases with time, the ratio $|\langle\hat{a}_{+}\hat{a}_{-}\rangle|^{2}/(n_{+}n_{-})$ decreases as the population grows, leading to a decrease of $g^{(2)}_{\pm}$. On the other hand, for hold times larger than 3 ms, the correlator begins to fall below 2 while the normalized number variance exceeds unity, a behavior not captured by Eq. (5) or its dissipative extension [54, 56]. In this regime, our entanglement witness is no longer satisfied and we cannot infer entanglement. The onset of this apparent loss of entanglement for the two modulation amplitudes occurs for similar mode populations.
As suggested in Ref. [57], this apparent loss of entanglement could indicate the onset of a late-time regime where higher order quasi-particle interactions become relevant and a richer phenomenology emerges, e.g. decoherence of the resonant modes [57], secondary spectral peaks as observed in similar experiments in hydrodynamic systems [58], or loss of Gaussianity [59, 60]. All these phenomena can be seen steps on the path towards thermalization for which we may be able to draw analogies to “reheating” in the cosmological era when the Universe transitions to a thermal state. Future work will investigate these effects.
Data availability – The data that support the findings of this manuscript are available in a public Zenodo repository at [61].
I Acknowledgments
@@
We acknowledge David Clément and Nicolas Pavloff for their careful reading of the manuscript and Maxime Jacquet for fruitful discussions. A. M. thanks all iTHEMS members, in particular Tsukasa Tada, for providing a supportive research environment. The research leading to these results has received funding from QuantERA Grant No. ANR-22-QUA2-000801 (MENTA) and ANR Grant No. 20-CE-47-0001-01 (COSQUA), the LabEx PALM (ANR-10-LABX-0039PALM). V.G. acknowledges the Région Ile-de-France in the framework of the DIM SIRTEQ program and the Quantum Saclay program FQPS (ANR-21-CMAQ-0002). R.D. acknowledges a PhD grant with reference 2024.03181.BD from the Portuguese Foundation for Science and Technology (FCT). S.R. is funded by the CNRS Chair in Physical Hydrodynamics (ANR-22-CPJ2-0039-01).
Appendix A End matter
A.1 Detector quantum efficiency
An accurate estimate of the quantum efficiency of the MCP detector is difficult. Earlier work on similar detectors on metastable helium yielded values ranging from 20% to 50% [64, 21, 65]. For plotting purposes, we take 25(5)%, compatible with the estimation made on our own detector. Since the entanglement witness depends on a normalized correlator, it depends very little on the exact value of the quantum efficiency. A higher quantum efficiency would shift the entanglement witness curve of Fig. 2 to the right but even a quantum efficiency of 100% would not change our conclusion on entanglement.
A.2 Auto-correlations
FIG. 4: (a) The auto-correlator computed with the same dataset as in Fig. 2, in which we vary the modulation depth for a constant hold time. Panels (b) and (c) show the corresponding auto-correlator for the dataset of Fig. 3 where the hold time is varied with a modulation depth of respectively 18% and 25%. The light square markers show the negative velocity sideband and the dark stars the positive velocity one. The error bars are computed from the expected error for a thermal law with the same detected atom number.
For purposes of comparison, we show in Fig. 4 the normalized two-body auto-correlators for the same data as that we used for Figs. 2 and 3. The normalized two-body auto-correlators are defined as
$g^{(2)}_{j}=\frac{G^{(2)}_{j}}{n_{j}^{2}}=\frac{\langle(\hat{a}_{j}^{\dagger})^{2}\hat{a}_{j}\rangle}{n_{j}^{2}}$ (6)
where $j=+$ or --. These auto-correlators are always consistent with the value 2 except in the case of very weak or very strong excitation. For weak excitation, the number of atoms in the sidebands is very low, resulting in large error bars, that we discuss below. In addition, the signal in the voxels is contaminated by other residual atoms. In the case of strong excitation, the decrease in the correlation may indicate a non-Gaussian state.
The error bars shown in Fig. 4 are not obtained using bootstrap techniques, but are instead estimated based on the number of repetitions and the mean detected atom number in the voxels. When the population is sufficiently large, this estimation yields results consistent with a bootstrap analysis. However, when the mean detected atom number is very low, the available statistics are insufficient to reliably estimate error bars. In some cases, for instance, no doublets are recorded at all, due to limited statistics. Thus, the error bars on $g^{(d)}$ are given by
$\Delta g^{(d)}=\frac{d!}{\sqrt{N_{\text{shot}}}}\sqrt{\frac{\text{Var}[(\hat{a}^{\dagger}_{j})^{d}\hat{a}_{j}^{d}]}{[(d!)n^{d}]^{2}}+d\frac{\text{Var}[\hat{a}^{\dagger}_{j}\hat{a}_{j}]}{n^{2}}}$ (7)
where the variance is evaluated using Eq. (11) only depends on the mean detected atom number and $N_{\text{shot}}$ is the number of experimental cycles. Here $d!$ represents the value of $g^{(d)}$ for a thermal statistics.
In the following, we derive Eq. (11) for a state with a thermal statistics. To do so, we need to evaluate
$\text{Var}[(\hat{a}^{\dagger})^{d}\hat{a}^{d}]=\braket{(\hat{a}^{\dagger})^{d}\hat{a}^{d}(\hat{a}^{\dagger})^{d}\hat{a}^{d}}-\braket{(\hat{a}^{\dagger})^{d}\hat{a}^{d}}^{2}$ (8)
where we drop the label $j$ to lighten notations. The second term in Eq. (8) is the $d$-body auto-correlator which is given by $G^{(d)}=\braket{(\hat{a}^{\dagger})^{d}\hat{a}^{d}}=(d!)\,n^{d}$ in the case of a thermal statistics [66]. To evaluate the first term, we normal order the creation and annihilation operators as [67]
$\begin{split}&(\hat{a}^{\dagger})^{d}\hat{a}^{d}(\hat{a}^{\dagger})^{d}\hat{a}^{d}=\ :\!\!\underbrace{(\hat{a}^{\dagger})^{d}\hat{a}^{d}(\hat{a}^{\dagger})^{d}\hat{a}^{d}}_{\text{no pair removed}}\!\!:\\ &\quad+:\underbrace{(\hat{a}^{\dagger})^{d}\hat{a}^{d-1}(\hat{a}^{\dagger})^{d-1}\hat{a}^{d}}_{\text{remove 1 pair}}:\times\underbrace{\binom{d}{1}\times\binom{d}{1}}_{\text{\# of pairs}}\\ &\quad+:\underbrace{(\hat{a}^{\dagger})^{d}\hat{a}^{d-2}(\hat{a}^{\dagger})^{d-2}\hat{a}^{d}}_{\text{remove 2 pairs}}:\times\underbrace{\binom{d}{2}\times\binom{d}{2}\times 2}_{\text{choose }\hat{a}\text{ then }\hat{a}^{\dagger}\text{, pair them}}\\ &\quad+:\underbrace{(\hat{a}^{\dagger})^{d}\hat{a}^{d-3}(\hat{a}^{\dagger})^{d-3}\hat{a}^{d}}_{\text{remove 3 pairs}}:\times\underbrace{\binom{d}{3}\times\binom{d}{3}\times 3\times 2}_{\text{choose }\hat{a}\text{ then }\hat{a}^{\dagger}\text{, pair them}}+...\end{split}$ (9)
from which we obtain
$\braket{(\hat{a}^{\dagger})^{d}\hat{a}^{d}(\hat{a}^{\dagger})^{d}\hat{a}^{d}}=\sum_{k=0}^{d}k!\binom{d}{k}^{2}G^{(2d-k)}.$ (10)
We therefore have
$\text{Var}[(\hat{a}^{\dagger})^{d}\hat{a}^{d}]=\sum_{k=0}^{d}k!\binom{d}{k}^{2}(2d-k)!n^{2d-k}-\left[d!n^{d}\right]^{2}.$ (11)

### v2

@@ Abstract
FIG. 3: Mean detected population (a) and cross-correlation (b) as a function of the hold time $\Delta t$. Data is shown for two different modulation depths (orange circles 18%, green squares 25%). The inset shows the normalized variance. At late times, entanglement cannot be inferred. Error bars denote one standard deviation computed using a bootstrap analysis.
[…]
Conclusion—We have observed entanglement between parametrically excited quasiparticle modes of a quantum fluid. This entanglement demonstrates the role of quantum fluctuations in seeding the parametric growth, in analogy with cosmological preheating. While it has its limitations [25], the relevance of the analogy is strengthened by the late-time dynamics. As noted above, the growing quasiparticle population leads to an interaction-dominated regime, of which the apparent loss of entanglement in Fig. 3 may be an indicator. This regime is characterized by a rich phenomenology, including decoherence of the resonant modes [61], secondary peaks as observed in similar hydrodynamic experiments [62], and loss of Gaussianity [63, 64]. These phenomena can be seen as steps towards thermalization, and taken together they are analogous to cosmological reheating [65]. In addition, as the BEC keeps breathing, the energy and coherence of the driving field will be lost due to backreaction from the produced quasiparticles, the effect of vacuum fluctuations playing again a crucial role [66]. Future work will investigate these effects.
Acknowledgements.
@@
Acknowledgments—We thank David Clément and Nicolas Pavloff for their careful reading of the manuscript and Maxime Jacquet for fruitful discussions. A. M. thanks Ryo Namba for discussions on preheating, and is grateful to all iTHEMS members, in particular Tsukasa Tada, for providing a supportive research environment. The research leading to these results has received funding from QuantERA Grant No. ANR-22-QUA2-000801 (MENTA), ANR Grant No. 20-CE-47-0001-01 (COSQUA), and the LabEx PALM (ANR-10-LABX-0039PALM). V.G. acknowledges funding from the Conseil Régional, Île-de-France (DIM SIRTEQ) and the Quantum Saclay program FQPS (ANR-21-CMAQ-0002). R.D. acknowledges a PhD grant ref. 2024.03181.BD from the Portuguese Foundation for Science and Technology (FCT). S.R. is funded by the CNRS Chair in Physical Hydrodynamics (ANR-22-CPJ2-0039-01).
Data availability—The data that support the findings of this manuscript are available in a public Zenodo repository at [67].